# Prioritizing Security Practice Adoption: Empirical Insights on Software Security Outcomes in the npm Ecosystem

### v1

@@ 3.1. Data Source
To collect data, we leveraged multiple sources. We selected the npm ecosystem for this study because npm is the largest open-source ecosystem, comprising 37 million package versions, with 74% of open-source codebases using JavaScript (ossra2024, ; sonatype2024, ). Scorecard analyzes packages with source repositories hosted on GitHub (openssf_scorecard_2023, ) and periodically evaluates popular open-source projects. Within Scorecard’s dataset of selected popular packages, 64% are npm packages. Additionally, our security outcome metrics require analyzing project dependencies, specifically counting vulnerabilities in dependencies and measuring dependency update patterns for MTTR and MTTU. To support our analysis, we collected the dependency graph of npm packages and mapped packages to their corresponding GitHub repositories using deps.dev (OSI, ). Deps.dev, developed and hosted by Google, provides comprehensive insights into OSS, including source code locations, dependency graphs, package metadata, licenses, releases, and vulnerabilities. Another platform used in our study is OSV (Open Source Vulnerabilities) (osv_adv, ). The OSV platform is an aggregator of vulnerabilities from different security advisories. Scorecard and Deps.dev depend on OSV to collect vulnerability information for projects and their dependencies. We also used the Libraries.io platform (LibrariesIO, ), which collects publicly available OSS information scraped from the internet, including GitHub.
3.2. Package Inclusion and Exclusion Criteria
@@
  • Initial Dataset: We began with a total of 2,119,044 unique npm packages extracted from the Deps.dev BigQuery dataset (depsdevBigQuery, ). This dataset contains the mapping between npm packages to GitHub repositories and captures all dependency edges within the npm.
  • Packages with GitHub Repositories: Of these, 1,241,732 packages were mapped to GitHub repositories. Since we must map each package to a Scorecard score via a GitHub repo, we exclude 877,312 packages that lack repository information, either due to hosting outside GitHub or missing repository URLs.
@@
  • Packages with Dependencies and Dependents: Further filtering identified 1,159,891 packages with at least one declared dependency, ensuring relevance to our security outcome metrics. Additionally, we are primarily interested in packages that practitioners may decide to use; hence, we refined the dataset further to include only packages that had at least one dependent, reducing the dataset to 264,413 packages.
@@
  • Ensuring Unique Repositories: Scorecard generates metric scores based on security practice adoption at the repository level rather than for individual packages. While we collected vulnerability data at the repository level (Section 3.4), MTTR and MTTU can only be measured at the package level. Consequently, if a repository contains multiple packages, all packages share the same Scorecard score and vulnerability data but may have different MTTR and MTTU values. Such inconsistency leads to duplicate rows with identical Scorecard metrics but differing security outcome metrics, which could mislead our machine learning model. To resolve this issue, we filtered the dataset to include only repositories that map to a single package. After applying this filtering step, we retained 172,999 unique repositories.
3.3. OpenSSF Scorecard Data
@@
Once we identified our set of 172,999 unique npm GitHub repositories, we collected OpenSSF Scorecard scores for each repository associated with a package. Our primary data source was the OpenSSF Scorecard dataset in Google BigQuery (ScorecardBigQuery, ). However, Scorecard only evaluates a subset of repositories, specifically popular packages and those explicitly requested by practitioners for inclusion. From Google BigQuery, we extracted Scorecard data for 761,602 repositories. However, not all of these repositories met our inclusion criteria mentioned in Section 3.2. For example, some lacked dependencies or dependents, while others had multiple packages mapped to single repositories, which was outside the scope of our study. Hence, many repositories in our dataset were not included in the original Scorecard dataset. To address these gaps, we ran Scorecard locally for any repositories missing from the BigQuery dataset, using the same configuration applied to populate BigQuery. After merging data from Google BigQuery and our local runs, we obtained Scorecard scores for 146,951 out of 172,999 repositories. The remaining repositories did not generate a score using Scorecard. Upon manual verification of 100 repositories, we found that Scorecard could not produce scores primarily because the repositories were either private or had been removed from GitHub.
Metrics inclusion and exclusion In this study, we studied 13 out of 18 Scorecard metrics, excluding CI-Tests, Contributors, Dependency-Update-Tool, Signed Releases, and Packaging, due to computational constraints or a lack of sufficient data. The CI-Tests, Contributors, and Dependency-Update-Tool metrics are not included in the Google BigQuery dataset due to the computational resources required to run these checks, which would hinder periodic updates. The exclusion was confirmed by the OpenSSF team, who specified the configuration in the project’s source code  [^1] https://github.com/ossf/scorecard/blob/76a04bfe40c25746b2633127270a9219a98a37ad/cron/config/config.yaml#L48. Scorecard generates a -1 score when no conclusive evidence of a practice’s adoption exists, either due to insufficient repository data or internal runtime errors. Signed Releases and Packaging were excluded, as over 90% of the repositories received a -1 score (Table 1). Prior research (zahan2023openssf, ) found that practitioners primarily used package registries for releases rather than GitHub, and no improvements in Scorecard’s methodology addressing this limitation were observed. Thus, we removed five metrics from the analysis.
Dangerous-Workflow, Pinned-Dependencies, and Token-Permissions metrics have less than 75% repositories with a -1 score (Table 1). Prior research by Zahan et al. (zahan2023openssf, ) highlighted limitations in Scorecard’s tool, particularly for these metrics that are used to assess security configuration in GitHub workflows and dependency management practice (Table 1). Previously, Scorecard assigned a maximum score of 10 for these metrics, even when no GitHub workflows or dependency metadata were present, which could mislead practitioners about positive security practice adoption. Our manual analysis of 50 repositories per metric confirms that Scorecard has improved the detection of these practices. Scorecard assigns a -1 score when repositories do not have any GitHub workflows, token information, or dependency metadata. A score of 0 now signifies that the practice exists with security misconfiguration, such as excessive GitHub token permissions, dangerous workflow patterns, or all dependencies are unpinned. We included these three metrics in our study because a -1 score represents non-adoption, while a score of 0 indicates security misconfigurations.
3.4. Security Outcome Metrics Data
@@
We collected our security outcome metrics for 146,951 npm packages with Scorecard data. Here, we describe our outcome metrics data collection process.
@@
Vul_Count metric: The Scorecard vulnerability metrics measure whether the project has open, unfixed vulnerabilities in its codebase or its dependencies using the OSV (osv_adv, ) service. Initially, we used Deps.dev BigQuery (depsdevBigQuery, ) and the Scorecard tool to collect the total vulnerability count. Deps.dev (OSI, ) includes Scorecard metrics for OSS. In deps.dev, we were able to map vulnerability data for 76,651 packages, and 70,300 remaining data were collected using Scorecard. The scorecard normalizes the vulnerability score, where a score of 10 indicates no vulnerabilities, 9 corresponds to one vulnerability, and 0 represents more than nine vulnerabilities. To obtain the exact vulnerability count, we first denormalized the score. For cases where the vulnerability score was 0 (indicating more than nine vulnerabilities), we ran the vulnerability metric separately using Scorecard to retrieve the raw, unnormalized data.
@@
MTTR and MTTU metric: In this study, we adopt the algorithms proposed by Rahman et al. (rahman2024characterizing, ) for computing MTTR and MTTU. Following their methodology, we contacted Google deps.dev team and considered deps.dev as our primary data source to collect comprehensive package versions and dependency versions and their release information for our dataset. Additionally, we retrieved CVE data for selected packages and their direct dependencies from OSV (osv_adv, ). Note that while MTTU captures all dependency updates, MTTR specifically accounts for updates addressing vulnerable dependencies. While we successfully collected Vul_Count and MTTU data for 146,951 packages, MTTR data was only available for 22,083 packages (15% of total dataset). The limitation arises because MTTR penalizes packages that did not update their dependencies despite a known vulnerability being fixed in an upstream version.
3.5. Control Variable Metrics Data
@@
We collected control variable data from different sources. For example, GitHub contribution count and package size are collected from Libraries.io (LibrariesIO, ). Then we got the number of downloads of a package in the past 12 months, which we collected from the public npm API [^2] https://api.npmjs.org/downloads/point/{period}[/{package}]. We calculated the repository age (in years) using the GitHub REST API (GitHubAPI2024, ) by retrieving the repository creation date and computing the difference between the current date and the creation date.
@@ 4. Data Preprocess
Before constructing machine learning (ML) modeling to answer RQ1, we performed data preprocessing steps to handle missing values, ensure data integrity, and assess statistical properties. Section 4.1 covers data imputation techniques to manage missing values. Section 4.2 discusses correlation and multicollinearity in our dataset. Figure 1 highlights the overview of our research workflow.
Figure 1. Research Workflow
4.1. Data Imputation
Data imputation is the process of replacing missing values to maintain data integrity, prevent biases in ML models, and ensure statistical validity (schafer2002missing, ). Imputation techniques help mitigate data loss, improve model accuracy, and prevent misleading conclusions caused by missing data. Depending on the nature of the missing data, multiple imputation techniques exist, such as mean/median imputation, categorical encoding, and predictive modeling-based imputations. In our dataset, Branch-Protection, Code-Review, Binary-Artifacts, Fuzzing, SAST, Security-Policy, License, and CII-Best-Practices each had less than 1% of packages with a -1 score. Table 1 shows the exact counts of packages with -1 for these metrics. On the contrary, Dangerous-Workflow, Pinned-Dependencies, and Token-Permissions had greater than 72% of repositories with a -1 score (Table 1). Our manual analysis found that Scorecard assigned -1 in two cases: (1) when there was no conclusive evidence of security practice adoption, and (2) due to internal runtime errors (section 3.3). Since not all metrics within a repository had a -1 score, removing entire repositories would result in unnecessary data loss. To handle missing data effectively, we differentiated between cases where -1 indicated non-adoption (e.g., missing GitHub workflows, token permissions, or dependency management practices) and cases where -1 resulted from internal runtime errors. For metrics where -1 was due to internal errors, we applied median imputation to maintain data completeness while minimizing distortion. For Dangerous-Workflow, Pinned-Dependencies, and Token-Permissions, where -1 indicated non-adoption rather than missing data, we applied categorical encoding imputation to preserve the distinction. Note that a score of 0 reflects detected security misconfigurations.
@@ 4.2. Correlation Analysis
To evaluate dataset normality, we applied the Henze-Zirkler multivariate normality test(henze1990class, ) separately to the RQ1 and RQ2 datasets  (Figure 1). The RQ1 dataset includes aggregated Scorecard score, Control Variables, Vul_Count, MTTR, and MTTU, while RQ2 leverage 12 Scorecard Metrics instead of aggregated Scorecard score. Both datasets failed the normality test (p ¡ 0.05), indicating deviations from a normal distribution. To further assess skewness, we flagged features with —skewness— ¿ 1 for potential transformation (hatem2022normality, ). Based on this, we applied log(x + 1) transformation to highly skewed variables (e.g., GitHub contributors, repository size, DependencyCount, downloads, MTTU, and Vul_Count) to mitigate skewness. Post-transformation, Q-Q plots, histograms, and statistical tests confirmed improved distribution symmetry. Since residual normality is important for regression analysis, we conducted D’Agostino-Pearson tests (keskin2006comparison, ) on regression residuals. Results showed that residuals remained non-normally distributed (p ¡ 0.05).
Given the non-normal distribution of our dataset, we proceeded with the Spearman correlation (zar2005spearman, ), a nonparametric method suitable for capturing monotonic relationships in non-normal data. We conducted correlation analysis for predictor variables (security practice and control variables) and target variables (security outcome metrics). Figure  2 presents the Spearman correlation heatmap for predictor variables, revealing weak to moderate correlations. The highest correlation is observed between contributors and code review (0.61), suggesting that repositories with more contributors are more likely to implement code review practices. The correlation between target variables reveals a strong positive correlation (0.7) between MTTR and MTTU, aligning with our expectation of potential overlap (rahman2024characterizing, ). In contrast, the Vul_Count metric exhibits a weak negative correlation with MTTR (-0.32) and MTTU (-0.04). Hence, we confirmed that MTTR and MTTU cannot be used as target variables in the same model due to their strong correlation since including both would introduce redundancy, bias, and multicollinearity. Additionally, we computed the Variance Inflation Factor (VIF) (o2007caution, ) for all predictor variables to confirm multicollinearity. For VIF computation, we found only “downloads” metrics had a high VIF score (15). According to study (o2007caution, ), VIF thresholds should be interpreted in context rather than strictly applied as exclusion criteria. Since the ”download” metric serves as a control variable in our study and does not exhibit strong correlations (¿0.8) with any other metrics, the high VIF alone is not sufficient grounds for removal (o2007caution, ).
@@ 5. RQ1: Overall security posture and its security outcomes
In this section, we discuss our RQ1: Is higher adoption of security practices associated with measurable improvements in security outcomes? Section 5.1 discusses our methodology to answer our RQ1, and Section 5.2 discusses our result.
Table 2. Summary Statistics of Security Outcomes
Statistic Vul_Count MTTR MTTU
Count 146,951 22,543 146,951
Mean 16.4 1,084.7 634.3
Std Dev 31.3 897.3 877.4
Min 0 1 0
Median 1.0 940.0 222.2
Max 637.0 4,709.0 4,869.6
@@ 5.1. RQ1: Methodology
For our RQ1, we evaluate the relationship between npm package repositories’ aggregated Scorecard score and control variables as predictors and security outcome metrics as the target variable (Figure 1). Table 2 presents the summary statistics of security outcomes. To visualize the relationship between aggregated Scorecard scores and security outcomes, we first binned repositories by aggregated Scorecard scores and analyzed mean security outcomes within each bin (Figure 3). Next, we applied regression analysis to quantify the association of the aggregated Scorecard score on outcomes. Finally, we conducted a causal analysis to assess whether the observed relationships reflect a causal effect.

### v2

@@ 3.2.1. Data Sources
To collect data, we leveraged multiple sources. We selected the npm ecosystem for this study because npm is the largest open-source ecosystem, comprising 37 million package versions, with 74% of open-source codebases using JavaScript (ossra2024, ; sonatype2024, ). Scorecard analyzes packages with source repositories hosted on GitHub (openssf_scorecard_2023, ) and periodically evaluates popular open-source projects. Within Scorecard’s dataset of selected popular packages, 64% are npm packages. Additionally, our security outcome metrics require analyzing project dependencies, specifically counting vulnerabilities in dependencies and measuring outdated and vulnerable dependency update patterns for MTTR and MTTU. To support our analysis, we collected the dependency graph of npm packages and mapped packages to their corresponding GitHub repositories using deps.dev (OSI, ). Deps.dev, developed and hosted by Google, provides comprehensive insights into OSS, including source code locations, dependency graphs, package metadata, licenses, releases, and vulnerabilities. Another platform used in our study is Open Source Vulnerabilities (OSV) (osv_adv, ). The OSV platform aggregates vulnerabilities from various security advisories. Scorecard and Deps.dev depend on OSV to collect vulnerability information for projects and their dependencies. We also used the Libraries.io platform (LibrariesIO, ), which collects publicly available OSS metadata scraped from the internet, including GitHub.
3.2.2. Package Inclusion and Exclusion Criteria
@@
  • Initial Dataset: We began with a total of 2,119,044 unique npm packages extracted from the Deps.dev BigQuery dataset (depsdevBigQuery, ). This dataset contains the mapping between npm packages and GitHub repositories and captures all dependency edges within npm.
  • Packages with GitHub Repositories: Of these, 877,312 packages were excluded because they lacked valid GitHub repository information, either because they were hosted outside GitHub or had missing repository URLs. Mapping packages to GitHub repositories was necessary because Scorecard runs directly on GitHub repositories. After this filtering step, 1,241,732 packages remained that could be successfully mapped to GitHub.
@@
  • Packages with Dependencies and Dependents: Further filtering identified 1,159,891 packages with at least one declared dependency, ensuring relevance to our security outcome metrics. Additionally, we are primarily interested in packages that practitioners may use; hence, we refined the dataset further to include only packages with at least one dependent, reducing the dataset to 264,413 packages.
@@
  • Ensuring Unique Repositories: Scorecard generates metric scores based on security practice adoption at the repository level rather than for individual packages. While we collected vulnerability data at the repository level (Section 3.2.4), MTTR and MTTU can only be measured at the package level. Consequently, if a repository contains multiple packages, all packages share the same Scorecard score and vulnerability data but may have different MTTR and MTTU values. Such inconsistency leads to duplicate rows with identical Scorecard metrics but differing security outcome metrics, which could mislead our machine learning model. To resolve this issue, we filtered the dataset to include only repositories that map to a single package. After applying this filtering step, we retained 172,999 unique repositories.
3.2.3. OpenSSF Scorecard Data
@@
Once we identified the set of 172,999 unique npm GitHub repositories, we collected OpenSSF Scorecard scores for each repository associated with a package. Our primary data source was the OpenSSF Scorecard dataset in Google BigQuery (ScorecardBigQuery, ). However, Scorecard only evaluates a subset of repositories, specifically popular packages and those explicitly requested by practitioners for inclusion. Hence, many repositories in our dataset were not included in the original Scorecard dataset. To address these gaps, we ran Scorecard locally for any repositories missing from the BigQuery dataset, using the same configuration applied to populate BigQuery. After merging data from Google BigQuery and our local runs, we obtained Scorecard scores for 145,817 out of 172,999 repositories. The remaining repositories did not generate a score using Scorecard. Upon manual verification of a random sample of 100 repositories and their Scorecard output, we found that Scorecard failed to produce scores due to HTTP 404 errors. The primary reason for the lack of scores was that the repositories were no longer accessible via public GitHub URLs. The repositories were either private or had been removed from GitHub.
Metrics exclusion: We began by collecting all 18 metrics provided by Scorecard. However, due to limited adoption and Scorecard measuring limitations, we excluded Fuzzing, CII-Best-Practices, Signed-Releases, and Packaging from our analysis (Table 1). Fuzzing and CII Best Practices were excluded because  99.8% of repositories scored 0, indicating minimal adoption of Scorecard recommendations across ecosystems. Signed Releases and Packaging were excluded, as almost 98% of the repositories received a -1 score due to not having any signed release or a package release workflow in GitHub. Prior research (zahan2023openssf, ) found that practitioners primarily used package registries for releases rather than GitHub, and no improvements in Scorecard’s methodology addressing this limitation were observed. We also removed the vulnerability metric as a practice because we use Vul_Count as a dependent variable. Thus, we removed five metrics from the analysis.
3.2.4. Security Outcome Metrics Data
@@
We collected our security outcome metrics for 145,817 npm packages with Scorecard data. Here, we describe our outcome metrics data collection process.
@@
Vul_Count metric: The Scorecard Vulnerability metric measures whether the project has open, unfixed vulnerabilities in its codebase or its dependencies using the OSV (osv_adv, ) service. Initially, we used the Scorecard tool to collect the total vulnerability count. The Scorecard tool normalizes the vulnerability score, where a score of 10 indicates no vulnerabilities, 9 corresponds to one vulnerability, and 0 represents more than nine vulnerabilities. To obtain the exact vulnerability count, we first denormalized the score. For cases where the vulnerability score was 0 (indicating more than nine vulnerabilities), we ran the vulnerability metric separately using Scorecard to retrieve the raw, unnormalized data and used the total number of vulnerabilities in the model.
@@
MTTR and MTTU metric: In this study, we adopt the algorithms proposed by Rahman et al. (rahman2024characterizing, ) for computing MTTR and MTTU. Following their methodology, we contacted the Googlee deps.dev team and considered deps.dev as our primary data source to collect comprehensive package versions, dependency versions and their release information for our dataset. Additionally, we retrieved CVE data for selected packages and their direct dependencies from OSV (osv_adv, ). Note that while MTTU captures all dependency updates, MTTR specifically accounts for updates addressing vulnerable dependencies. MTTR needs a vulnerability in direct dependency, then the availability of a new fix version, and the adoption time to adopt the new fix version by the dependent. While we successfully collected Vul_Count and MTTU data for 145,817 packages, MTTR data was only available for 22,412 packages (15% of total dataset).
3.2.5. Control Variable Metrics Data
@@
We collected control variable data from different sources. GitHub contribution count and package size are collected from Libraries.io (LibrariesIO, ). Then we got the number of downloads of a package in the past 12 months, which we collected from the public npm API [^2] https://api.npmjs.org/downloads/point/{period}[/{package}]. We calculated the repository age (in years) and commit staleness dates using the GitHub REST API (GitHubAPI2024, ). The number of dependency information is collected from Deps.dev (depsdevBigQuery, ).
@@ 4. Data Preprocess
Before constructing the machine learning (ML) modeling to answer RQs, we performed data preprocessing steps to handle missing values, ensure data integrity, and assess statistical properties. Section 4.1 covers feature decomposition and imputation techniques to manage missing values. Section 4.2 discusses correlation and multicollinearity in our dataset.
4.1. Feature Decomposition and Imputation
Our manual analysis found that Scorecard assigned -1 in two cases: (1) when there was no conclusive evidence of security practice adoption; and (2) due to internal runtime errors. Since not all metrics within a repository had a -1 score (Table 1), excluding an entire repository due to a -1 in any one of the 18 metrics would result in unnecessary data loss. To handle missing data effectively, we implemented a missing data pattern decomposition strategy based on Scorecard score reasoning. For high-missing features CI-Tests, Pinned-Dependencies, Token-Permissions, and Dangerous-Workflow (each with $>$50% -1 scores), a score of -1 indicates that the repository lacks the relevant configuration. Here, -1 for CI-Tests indicates no pull requests to confirm CI workflows were present; for Pinned-Dependencies, it suggests the absence of declared dependencies in GitHub; for Token-Permissions, it reflects that no GitHub token settings were configured; and for Dangerous-Workflow, -1 means no GitHub Actions workflows were found to test workflow patterns. To preserve the distinction, we decomposed each feature into two components: a binary indicator of relevant configuration presence (1 for values 0-10, 0 for -1) and scores that capture adoption effectiveness (keeping original discrete values 1-10, with both -1 and 0 scores mapped to 0). For low-missing features ($<$1% -1 values), where -1 was due to internal errors of Scorecard, we applied median imputation (schafer2002missing, ) to maintain data completeness while minimizing distortion.
@@ 4.2. Correlation Analysis
To evaluate dataset normality, we applied the Henze-Zirkler multivariate normality test (henze1990class, ) separately to the RQ1 and RQ2 datasets. The RQ1 dataset includes aggregate scorecard scores, the control Variables, Vul_Count, MTTR, and MTTU. The RQ2 dataset contains 13 Scorecard practice metrics, control variables, and Vul_Count, MTTR, and MTTU. Both datasets failed the normality test (p $<$ 0.05), indicating deviations from a normal distribution. To further assess skewness, we flagged features with $|skewness|>$ 1 for potential transformation (hatem2022normality, ). Based on skewness, we applied a log(x + 1) transformation to highly skewed control variables (GitHub contributors, repository size, DependencyCount, downloads) to mitigate skewness. Post-transformation, Q-Q plots, histograms, and statistical tests confirmed improved distribution symmetry. We did not apply transformation to the practice or outcome metrics to preserve their original scale and interoperability. Since residual normality is important for regression analysis, we conducted D’Agostino-Pearson tests (keskin2006comparison, ) on regression residuals. Results showed that residuals remained non-normally distributed (p $<$ 0.05).
Given the non-normal distribution of our dataset, we proceeded with the Spearman correlation (zar2005spearman, ), a nonparametric method suitable for capturing monotonic relationships in non-normal data. We conducted a correlation analysis for predictor variables (security practice and control variables) and target variables (security outcome metrics). Figure  1 presents the Spearman correlation heatmap for predictor variables. The highest correlations ( 1) were observed among Dangerous-Workflow, Pinned-Dependencies, and Token-Permissions. To avoid multicollinearity and redundancy in the model (guyon2003introduction, ), we removed Dangerous-Workflow and Token-Permissions, retaining only Pinned-Dependencies as a representative feature. The correlation between target variables reveals a strong positive correlation (0.7) between MTTR and MTTU, aligning with our expectation of potential overlap (rahman2024characterizing, ). In contrast, the Vul_Count metric exhibits a weak negative correlation with MTTR (-0.3) and MTTU (-0.4). Hence, we confirmed that MTTR and MTTU cannot be used as target variables in the same model due to their strong correlation, since including both would introduce redundancy, bias, and multicollinearity. Additionally, we also computed the Variance Inflation Factor (VIF) (o2007caution, ) for all predictor variables to confirm multicollinearity. For the VIF computation, we did not find any feature that exhibited a VIF score greater than 5, indicating no multicollinearity.
@@ 5. RQ1: Overall security posture and its security outcomes
In this section, we discuss our RQ1: Is higher adoption of aggregated security practices associated with measurable improvements in security outcomes? Section 5.1 discusses the methodology of RQ1, and Section 5.2 discusses the results.
@@ 5.1. RQ1: Methodology
For RQ1, we evaluate the relationship between the aggregated Scorecard score of npm package repositories and control variables as predictors, and security outcome metrics as the target variable. To visualize the relationship between aggregated Scorecard scores and security outcomes, we first binned repositories by aggregated Scorecard scores and analyzed mean security outcomes within each bin (Figure 2). Next, we applied regression analysis to quantify the association of the aggregated Scorecard score with outcomes. However, regression analysis does not establish causality. We implemented a Propensity Score Matching (PSM) framework with caliper-based matching, covariate balance diagnostics, and segment-level analysis to estimate the causal effect of security practice adoption on outcome metrics, controlling for project characteristics.
5.1.1. Data Expolation
To visualize initial relationships, we binned repositories by aggregated Scorecard scores into five equal-width bins and calculated mean security outcomes within each bin. The binning approach provides an intuitive view of how security outcomes trend across different score ranges before applying statistical modeling.